OpenStudy (anonymous):

Help! How do I calculate the derivative of y: xy^4+x^2y=x+3y

OpenStudy (anonymous):

\[xy ^{4}+x^{2}y=x+3y\]

OpenStudy (anonymous):

y4 + 4xy^3 dy/dx + 2xy +x^2 dy/dx = 1+ 3 dy/dx

OpenStudy (anonymous):

I'd like to know how to do it. Can someone explain what to do with this sort of problem?

OpenStudy (anonymous):

do u know product rule..

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What am I supposed to do? I still don't know where to start.

OpenStudy (yttrium):

You know of implicit differentiation?

OpenStudy (yttrium):

Just derive them term by term. Let me give you an example, in the right hand side. The derivative of x+3y is simply 1+3y'

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

Oh, that's it?

OpenStudy (yttrium):

Not yet. Can you do the derivative in the left hand side?

OpenStudy (anonymous):

Yeah. Using the product rule?

OpenStudy (yttrium):

Yeah. You can show me the result of your answer for clarification. :)

OpenStudy (anonymous):

Alright, great! I'm working the problem out right now!

OpenStudy (yttrium):

any updates?

OpenStudy (anonymous):

I got 4xy^3+y^4+2x^3y=1+3y'

OpenStudy (yttrium):

are you sure?

OpenStudy (yttrium):

another thing: take note that derivative of y is y'

OpenStudy (anonymous):

Oh, whoops, I multiplied in a place where I should have added. Now I got 4xy'^3+y'^4+2x^y'+x^2=1+3y'

OpenStudy (yttrium):

Still no. It should be: 4xy^3 y' + y^4 +2xy +x^2 y' = 1+3y'

OpenStudy (anonymous):

Oh okay. I made some more typos in my last answer, whoops! but is that all there is? That's it?

OpenStudy (yttrium):

Not yet of course. You need to isolate y' and equate the remaining terms to it.